# LaCATODA 2026: The 10th Linguistic and Cognitive Approaches to Dialog Agents Workshop at the 40th AAAI conference, Janurary 20–27, 2026, Singapore Expo

### v2

@@ 1 Introduction
Large Language Models (LLMs) struggle with domain-specific multi-turn dialogues despite Retrieval-Augmented Generation (RAG) enhancements [1]. Customer service dialogues present unique challenges: temporal dependencies creating complex reference chains across turns, and dynamically evolving conversation objectives that traditional RAG approaches fail to address effectively.
Current research follows two separate trajectories: Conversation RAG methodologies [2, 3] optimize dialogue history representation through semantic matching, while GraphRAG approaches [4, 5, 6] leverage knowledge graphs for reasoning over static domain knowledge. Both have fundamental limitations—Conversation RAG lacks structural guidance for conversation flow, while GraphRAG fails to model dynamic intent transitions essential for coherent multi-turn dialogues.
@@ 2 Related Works
RAG enhances Large Language Models (LLMs) by integrating external knowledge sources, improving response accuracy and informativeness in dialogue systems [1]. Recent advancements include memory-based approaches like DH-RAG [7] and SELF-multi-RAG [8], evaluation frameworks like MTRAG [9], and specialized applications combining knowledge graphs for customer service [10] and real-time problem detection with FAQ retrieval [11].
GraphRAG enhances traditional RAG by integrating knowledge graphs for structured semantic retrieval and multi-hop reasoning [4], better capturing contextual dependencies in multi-turn dialogues [5]. Key advantages include enhanced retrieval accuracy through preserved structural relationships [10], effective processing of multi-hop queries requiring cross-fragment reasoning [4, 6], context continuity maintenance via entity relationship tracking [10], and improved answer interpretability through visualized reasoning paths [12].
Conversation RAG addresses multi-turn dialogue challenges in customer service, where user queries often depend on historical interactions [2, 3].Key advancements include context-aware frameworks like DH-RAG [7] and SELF-multi-RAG [8]; structured retrieval approaches such as KG-RAG [10]; query decomposition techniques like Collab-RA [13]; and specialized evaluation benchmarks including RAD-Bench [14], MTRAG [9], and ChatQA [15]. While these approaches have advanced independently, no unified solution simultaneously addresses both contextual coherence and structured reasoning challenges.
@@ 4.1.2 Evaluation Metrics
LLM-as-Judge Evaluation: We utilized Claude 3.7 Sonnet [16] as an independent evaluator to assess both retrieved response quality and generated responses. For each evaluation instance, the LLM was provided with the complete dialogue history, the current user query, anonymized system responses generated by different systems, and the ground truth response for reference. For each comparison, LLM compared outputs from all systems across five dimensions (relevance, helpfulness, style consistency, contextual appropriateness, and professionalism) and provided numerical ratings (1-10) for each system, with the highest-rated system designated as "winner" for that example. For retrieved quality evaluation, we excluded the "professionalism" dimension, and removed ground truth responses in the LLM prompt. The complete prompt template for reponse generation is provided in Appendix E.

### v3

@@ 1 Introduction
Large Language Models (LLMs) struggle with domain-specific multi-turn dialogues despite Retrieval-Augmented Generation (RAG) enhancements [Lewis2020]. Customer service dialogues present unique challenges: temporal dependencies creating complex reference chains across turns, and dynamically evolving conversation objectives that traditional RAG approaches fail to address effectively.
Current research follows two separate trajectories: Conversation RAG methodologies [researchgate2025learning, ye2024boosting] optimize dialogue history representation through semantic matching, while GraphRAG approaches [graphrag2025, graphrag2024, zhang2025credible] leverage knowledge graphs for reasoning over static domain knowledge. Both have fundamental limitations—Conversation RAG lacks structural guidance for conversation flow, while GraphRAG fails to model dynamic intent transitions essential for coherent multi-turn dialogues.
@@ 2 Related Works
RAG enhances Large Language Models (LLMs) by integrating external knowledge sources, improving response accuracy and informativeness in dialogue systems [Lewis2020]. Recent advancements include memory-based approaches like DH-RAG [zhang2025dhrag] and SELF-multi-RAG [self-rag], evaluation frameworks like MTRAG [katsis2025mtrag], and specialized applications combining knowledge graphs for customer service [xu2024retrieval] and real-time problem detection with FAQ retrieval [agrawal2024beyond].
GraphRAG enhances traditional RAG by integrating knowledge graphs for structured semantic retrieval and multi-hop reasoning [graphrag2025], better capturing contextual dependencies in multi-turn dialogues [graphrag2024]. Key advantages include enhanced retrieval accuracy through preserved structural relationships [xu2024retrieval], effective processing of multi-hop queries requiring cross-fragment reasoning [graphrag2025, zhang2025credible], context continuity maintenance via entity relationship tracking [xu2024retrieval], and improved answer interpretability through visualized reasoning paths [linders2025knowledge].
Conversation RAG addresses multi-turn dialogue challenges in customer service, where user queries often depend on historical interactions [researchgate2025learning, ye2024boosting].Key advancements include context-aware frameworks like DH-RAG [zhang2025dhrag] and SELF-multi-RAG [self-rag]; structured retrieval approaches such as KG-RAG [xu2024retrieval]; query decomposition techniques like Collab-RA [collab-rag]; and specialized evaluation benchmarks including RAD-Bench [kuo2024radbench], MTRAG [katsis2025mtrag], and ChatQA [chatqa]. While these approaches have advanced independently, no unified solution simultaneously addresses both contextual coherence and structured reasoning challenges.
@@ 4.1.2 Evaluation Metrics
LLM-as-Judge Evaluation: We utilized Claude 3.7 Sonnet [sonnet_aws] as an independent evaluator to assess both retrieved response quality and generated responses. For each evaluation instance, the LLM was provided with the complete dialogue history, the current user query, anonymized system responses generated by different systems, and the ground truth response for reference. For each comparison, LLM compared outputs from all systems across five dimensions (relevance, helpfulness, style consistency, contextual appropriateness, and professionalism) and provided numerical ratings (1-10) for each system, with the highest-rated system designated as "winner" for that example. For retrieved quality evaluation, we excluded the "professionalism" dimension, and removed ground truth responses in the LLM prompt. The complete prompt template for reponse generation is provided in Appendix E.